# Offline Learning for Combinatorial Multi-armed Bandits

### v1

@@ 4.3 Offline Learning for Influence Maximization with Extension to Node-level Feedback
Influence maximization (IM) is the task of selecting a small number of seed nodes in a social network to maximize the influence spread from these nodes, which has been applied in various important applications such as viral marketing, epidemic control, and political campaigning [96, 97, 98]. IM has been intensively studied over the past two decades under various diffusion models, such as the independent cascade (IC) model [97], the linear threshold (LT) model [117], and the voter model [118]], as well as different feedback such as edge-level and node-level feedback models [106]. For the edge-level feedback model, IM smoothly fits into our framework by viewing each edge as the base arm, which can obtain the theoretical result similar to our previous two applications. In this section, we consider a more realistic yet challenging setting where we can only obverse the node-level feedback, showing that our framework still applies as long as we can construct a high probability lower bound (LCB) for each base arm (edge).
@@
Similar to [107], we assume these $n$ influence cascades are generated independently from a seed set distribution $S_{t,0}\sim\mathbb{D}_{\mathcal{S}}$, and given $S_{t,0}$, the cascades are generated according to the IC diffusion process. For each node $v\in\mathcal{V}$, we use $q_{v}=\Pr\left[v\in S_{t,0}\right]$ to denote the probability that the node $v$ is selected by the experimenter in the seed set $S_{t,0}$. We use $p_{G}(\bar{v})\vcentcolon=\Pr\left[v\notin S_{t,1}\right]$ to denote the probability that the node $v$ is not activated in one time step when the graph is $G$. For any two nodes $u,v\in\mathcal{V}$, we use $p_{G}(\bar{v}|u)\vcentcolon=\Pr\left[v\notin S_{t,1}|u\in S_{t,0}\right]$ and $p_{G}(\bar{v}|\bar{u})\vcentcolon=\Pr\left[v\notin S_{t,1}|u\notin S_{t,0}\right]$ to denote the probability that the node $v$ is not activated in one time step conditioned on whether the node $u$ is in the seed set $S_{t,0}$ or not, respectively. We also assume $\mathbb{D}_{\mathcal{S}}$ is a product distribution, i.e., each node $u\in\mathcal{V}$ is selected as a seed node in $S_{t,0}$ independently. Similar to [107], we also need an additional 1.
@@
Inspired by [107], for each node-level feedback data $t\in[n]$, we only use the seed set $S_{t,0}$ and the active nodes in the first diffusion step $S_{t,1}$ to construct the LCB.
@@
To find out an action $\widehat{S}$ such that $\sigma(\widehat{S};G)\geq(\alpha-\epsilon)\sigma(S^{*};G)$, our algorithm requires that $n\geq\tilde{O}\left(\frac{V^{2}d_{\max}^{2}}{\epsilon^{2}\eta\gamma^{3}}\right)$, which improves the existing result by at least a factor of $\tilde{O}\left(\frac{V^{4}}{k^{2}d^{2}_{\max}\eta}\right)$, owing to our variance-adaptive LCB construction and the tight CMAB-T analysis. We also relax the assumption regarding 1, where we require bounded $q_{u},p(\bar{v})$ only for $(u,v)\in\mathcal{E}(s^{*})$, since we use LCB $\underaccent{\bar}{p}_{uv}$. Chen et al. [107], instead, needs bounded $q_{u},p(\bar{v})$ for all $(u,v)\in\mathcal{E}$ as they directly use the empirical mean of $p_{uv}$.
5 Experiments
(a) Synthetic Dataset
(b) Real-world Dataset
Figure 1: Suboptimality gaps for cascading bandit application.
Figure 2: Comparison of different offline data generation methods.
(a) Synthetic Dataset
(b) Real-world Dataset
Figure 3: Suboptimality gaps for LLM cache application.
We now present experimental results on both synthetic and real-world datasets. For cascading bandits on the application of learning to rank, in the synthetic setting, item parameters $\mu_{i}$ are drawn from $U[0,1]$, and in each round $t$, a ranked list $S_{t}$ of $K$ items is randomly sampled. Fig. 1(a) shows that compared to CUCB-Offline [8], which is an offline adaptation of CUCB for our setting, and EMP [119], CLCB (Algorithm 1) reduces suboptimality gaps by 47.75% and 20.02%, respectively. For real-world evaluation, we use the Yelp datasethttps://www.yelp.com/dataset, where users rate businesses. We randomly select $m=200$ rated items per user and recommend up to $K$ items to maximize the probability of user engagement. The unknown probability $\mu_{i}$ is derived from Yelp, and cascading feedback is collected. Fig. 1(b) compares suboptimality gaps over $n=100$ rounds for $K=4,8$ , with a logarithmic scale on the y-axis. Note that as $K$ increases, the expected reward also changes, thus reducing suboptimality gaps. CLCB consistently achieves the lowest suboptimality gap.
Similarly, we conduct experiments in the LLM cache setting. In the synthetic setup, we simulate 100 distinct queries with a cache size of 40, following a power-law frequency distribution ($\alpha=0.9$) as in [34]. As shown in Fig. 3(a), our CLCB-LLM-C algorithm outperforms LFU [34] and LEC [34], achieving at least 1.32$\times$ improvement. For real-world evaluation, we use the SciQ dataset [120]. We evaluate GPT-4-o with the “o200k_base” encoding with cache sizes $K=10$ and $K=20$, where cost is defined by OpenAI’s API pricing with the tiktoken library [121]. Fig. 3(b) shows that CLCB-LLM-C (Algorithm 3) reduces costs by up to 36.01% and 20.70%, compared to LFU and LEC. Moreover, a larger $K$ shows a lower suboptimality gap, which is consistent with Theorem 3. Further details on experimental setups, results, and additional comparisons can be found in Appendix G.
The appendix is organized as follows.
  • In Appendix A, we prove the upper bound of the suboptimal gap
    • under the infinity-norm TPM data coverage condition (3)
    • under 1-norm TPM data coverage condition (4).
  • In Appendix B, we prove
    • the lower bound of suboptimal gap for the $k$-path problem with semi-bandit feedback.
  • In Appendix C, we prove
    • the gap upper bound for the offline learning problem in cascading bandits.
  • In Appendix D, we prove
    • the standard gap upper bound of for offline learning in LLM cache
    • the improved gap upper bound for offline learning in LLM cache
    • the improved regret upper bound for online streaming LLM cache
  • In Appendix E, we prove
    • the gap upper bound for the influence maximization under the node-level feedback.
  • In Appendix F, we prove auxiliary lemmas that serve as important ingredients for our analysis.
  • In Appendix G, we provide the additional experimental results.
Appendix A Proof for the Upper Bound Result

### v2

@@ 4.3 Offline Learning for Influence Maximization with Extension to Node-level Feedback
Influence maximization (IM) is the task of selecting a small number of seed nodes $S$ in a social network $G(\mathcal{V},\mathcal{E},p)$ to maximize the influence spread $\sigma(S;G)$ from these nodes. IM has been intensively studied over the past two decades under various diffusion models, such as the independent cascade (IC) model (Kempe et al., 2003b), the linear threshold (LT) model (Chen et al., 2010), as well as different feedback such as edge-level and node-level feedback models. For the edge-level feedback model, IM smoothly fits into our framework by viewing each edge weight $p_{uv}$ for $(u,v)\in\mathcal{E}$ as the base arm. In this section, we consider a more realistic yet challenging setting where we can only obverse the node-level feedback (Chen et al., 2020, 2021), showing that our framework still applies as long as we can construct a high probability lower bound (LCB) for each base arm (edge). Due to space constraints, we present only our main result here. The detailed setting, algorithm, and theoretical analysis can be found in Appendix C.
Theorem 4.
Under 1, suppose the number of data $n\geq\frac{392\log(\frac{12nE}{\delta})}{\eta\cdot\gamma}$. Letting $\hat{S}$ be the seed set returned by Algorithm 3, then it holds with probability at least $1-\delta$,
$\alpha\sigma(S^{*};G)-\sigma\left(\hat{S};G\right)$ (12)
$\leq 48\sqrt{6}\sqrt{\frac{V^{2}d^{2}_{\max}\sigma^{2}(S^{*};G)\cdot\log(\frac{12nE}{\delta})}{\eta\cdot\gamma^{3}\cdot n}},$
where $d_{\max}$ is the maximum out-degree, $\gamma,\eta$ are lower bounds for seed sampling probability and the activation probability, respectively, as in 1.
Remark 6 (Discussion).
To find out an action $\widehat{S}$ such that $\sigma(\widehat{S};G)\geq(\alpha-\epsilon)\sigma(S^{*};G)$, our algorithm requires that $n\geq\tilde{O}\left(\frac{V^{2}d_{\max}^{2}}{\epsilon^{2}\eta\gamma^{3}}\right)$, which improves the existing result by a factor of $\tilde{O}\left(\frac{V^{4}}{k^{2}d^{2}_{\max}\eta}\right)$, owing to our variance-adaptive LCB construction and the tight CMAB-T analysis. We also relax the assumption of Chen et al. (2021) for the seed sampling probability and activation probability, owing to the usage of LCBs, rather than directly using the empirical mean of edge weight $p_{uv}$. See Appendix C for the detailed discussion.
5 Experiments
(a) Synthetic Dataset
(b) Real-world Dataset
Figure 1: Suboptimality gaps for cascading bandit application.
Figure 2: Comparison of different offline data generation methods.
(a) Synthetic Dataset
(b) Real-world Dataset
Figure 3: Suboptimality gaps for LLM cache application.
We now present experimental results on both synthetic and real-world datasets. Each experiment was conducted over 20 independent trials. For cascading bandits on the application of learning to rank, in the synthetic setting, item parameters $\mu_{i}$ are drawn from $U[0,1]$, and in each round $t$, a ranked list $S_{t}$ of $K$ items is randomly sampled. Fig. 1(a) shows that compared to CUCB-Offline (Chen et al., 2016), which is an offline adaptation of CUCB for our setting, and EMP (Liu et al., 2021), which always selects the action based on the empirical mean of rewards, CLCB (Algorithm 1) reduces suboptimality gaps by 47.75% and 20.02%, respectively. For real-world evaluation, we use the Yelp dataset [^3] https://www.yelp.com/dataset, where users rate businesses (Dai et al., 2024c). We randomly select $m=200$ rated items per user and recommend up to $K$ items to maximize the probability of user engagement. The unknown probability $\mu_{i}$ is derived from Yelp, and cascading feedback is collected. Fig. 1(b) compares suboptimality gaps over $n=100$ rounds for $K=4,8$ , with a logarithmic scale on the y-axis. Note that as $K$ increases, the expected reward also changes, thus reducing suboptimality gaps. CLCB consistently achieves the lowest suboptimality gap.
Similarly, we conduct experiments in the LLM cache setting. In the synthetic setup, we simulate 100 distinct queries with a cache size of 40, following a power-law frequency distribution ($\alpha=0.9$) as in (Zhu et al., 2023). As shown in Fig. 3(a), our CLCB-LLM-C algorithm outperforms LFU (which evicts the least frequently accessed items to optimize cache usage) (Zhu et al., 2023) and LEC (which minimizes inference cost by evicting items with the lowest estimated expected cost) (Zhu et al., 2023), achieving at least 1.32$\times$ improvement. For real-world evaluation, we use the SciQ dataset (Welbl et al., 2017). We evaluate GPT-4-o with the “o200k_base” encoding with cache sizes $K=10$ and $K=20$, where cost is defined by OpenAI’s API pricing with the tiktoken library (OpenAI, 2025). Fig. 3(b) shows that CLCB-LLM-C (Algorithm 2) reduces costs by up to 36.01% and 20.70%, compared to LFU and LEC. Moreover, a larger $K$ shows a lower suboptimality gap, which is consistent with Theorem 3. Further details on experimental setups, results, and additional comparisons can be found in Appendix J.
6 Conclusion and Future Directions
In this paper, we introduce Off-CMAB, the first offline learning framework for CMAB. We propose two novel data coverage conditions and develop a provably sample-efficient CLCB algorithm, matching the lower bound up to logarithmic factors. We show the practical usefulness of our framework via three diverse applications—learning to rank, LLM caching, and social influence maximization—achieving new or improved theoretical results. These results are further validated through extensive experiments on both synthetic and real-world datasets. Looking ahead, an exciting direction is to develop variance-adaptive algorithms to further improve our theoretical guarantees. Additionally, extending our framework to offline RL with combinatorial action spaces is another promising direction for future research.
Acknowledgement
The work is supported by NSF CNS-2103024 and the Office of Naval Research under grant N000142412073. The work of John C.S. Lui was supported in part by the RGC GRF-14202923. The work of Jinhang Zuo was supported by CityUHK 9610706.
Impact Statement
This paper presents a theoretical study on multi-armed bandits and reinforcement learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Extended Related Works
A.1 Combinatorial Multi-armed Bandits
The combinatorial multi-armed bandit (CMAB) problem has been extensively studied over the past decade, covering domains such as stochastic CMAB (Gai et al., 2012; Kveton et al., 2015c; Combes et al., 2015; Chen et al., 2016; Wang & Chen, 2017; Merlis & Mannor, 2019; Saha & Gopalan, 2019; Agrawal et al., 2019; Liu et al., 2022, 2024b), adversarial CMAB (György et al., 2007; Uchiya et al., 2010; Cesa-Bianchi & Lugosi, 2012; Bubeck et al., 2012; Audibert et al., 2014; Neu, 2015; Han et al., 2021), and hybrid best-of-both-worlds settings (Zimmert et al., 2019; Ito, 2021; Tsuchiya et al., 2023). Contextual extensions with linear or nonlinear function approximation have also been explored (Qin et al., 2014; Takemura et al., 2021; Liu et al., 2023a; Chen et al., 2018; Nika et al., 2020; Choi et al., 2024; Hwang et al., 2023; Liu et al., 2024b).
Our work falls within the stochastic CMAB with semi-bandit feedback domain, first introduced by Gai et al. (2012), with a specific focus on CMAB with probabilistically triggered arms (CMAB-T). Chen et al. (2016) introduced the concept of arm triggering processes for applications like cascading bandits and influence maximization, proposing the CUCB algorithm with a regret bound of $O(B_{1}\sqrt{mKT\log T}/p_{\min})$ regret bound under the 1-norm smoothness condition with coefficient $B_{1}$. Subsequently, Wang & Chen (2017) refined this result, proposed a stronger 1-norm triggering probability modulated (TPM) $B_{1}$ smoothness condition, and employed triggering group analysis to eliminate the $1/p_{\min}$ factor from the previous regret bound. More recently, Liu et al. (2022) leveraged the variance-adaptive principle to propose the BCUCB-T algorithm, which further reduces the regret’s dependency on action-size from $O(K)$ to $O(\log K)$ under the new variance and triggering probability modulated (TPVM) condition. While inspired by these works, our study diverges by addressing the offline CMAB setting, where online exploration is unavailable, and the focus is on minimizing the suboptimality gap rather than regret.
Another notable line of work considers CMAB with full bandit feedback (György et al., 2007; Cesa-Bianchi & Lugosi, 2012; Niazadeh et al., 2021; Fourati et al., 2023; Nie et al., 2023; Fourati et al., 2024b, a; Sun et al., 2025). In their setting, the feedback only provides aggregate rewards for the entire super arm, often resulting in higher regret (e.g., $O(T^{2/3})$) and requiring fundamentally different oracle designs. In contrast, our setting (CMAB with semi-bandit feedback) assumes semi-bandit feedback, where the learner observes individual arm-level feedback for selected arms (i.e., components of the super arm). This enables more informative learning and allows us to construct accurate base-arm estimators for use in our oracles, leading to an $O(T^{1/2})$ regret bound. Moreover, prior full-bandit approaches often rely on additional structural assumptions such as submodularity to achieve these bounds. Similarly, our approach leverages smoothness assumptions on the reward function to ensure statistical efficiency in the offline regime.
A.2 Offline Bandit and Reinforcement Learning
Offline reinforcement learning (RL), also known as “batch RL”, focuses on learning from pre-collected datasets to make sequential decisions without online exploration. Initially studied in the early 2000s (Ernst et al., 2005; Riedmiller, 2005; Lange et al., 2012), offline RL has gained renewed interest in recent years (Levine et al., 2020).
From an empirical standpoint, offline RL has achieved impressive results across diverse domains, including robotics (Singh et al., 2021), healthcare (Liu et al., 2020), recommendation systems (Chen et al., 2023), autonomous driving (Kiran et al., 2020), and large language model fine-tuning and alignment (Casper et al., 2023). Algorithmically, offline RL approaches can be broadly categorized into policy constraint methods (Fujimoto et al., 2018; Kumar et al., 2019), pessimistic value/policy regularization (Haarnoja et al., 2018; Kumar et al., 2020), uncertainty estimation (Agarwal et al., 2019), importance sampling (Jiang & Li, 2015; Nachum et al., 2019), imitation learning (Fujimoto & Gu, 2021; Chen et al., 2019), and model-based methods (Kidambi et al., 2020; Yu et al., 2021).
Theoretically, early offline RL studies relied on strong uniform data coverage assumptions (Szepesvari & Munos, 2005; Chen & Jiang, 2019a; Wang et al., 2019; Xie et al., 2021a). Recent works have relaxed these assumptions to partial coverage for tabular Markov Decision Processes (MDPs) (Rashidinejad et al., 2021; Yin et al., 2021; Shi et al., 2022; Li et al., 2022b), linear MDPs (Jin et al., 2020b; Chang et al., 2021; Bai et al., 2022), and general function approximation settings (Rashidinejad et al., 2022; Zanette et al., 2021; Xie et al., 2021b; Zanette & Wainwright, 2022).
Offline bandit learning has also been explored in multi-armed bandits (MAB) (Rashidinejad et al., 2021), contextual MABs (Rashidinejad et al., 2021; Jin et al., 2020b; Li et al., 2022a), and neural contextual bandits (Nguyen-Tang et al., 2021b, a).
While our work leverages the pessimism principle and focuses on partial coverage settings, none of the aforementioned offline bandit or RL studies address the combinatorial action space, which is the central focus of our work. Conversely, recent work in the CMAB framework demonstrates that episodic tabular RL can be viewed as a special case of CMAB (Liu et al., 2024b). Building on this connection, our proposed framework can potentially extend to certain offline RL problems, offering a unified approach to tackle both combinatorial action spaces and offline learning.
A.3 Related Offline Learning Applications.
Cascading bandits, a classical online learning-to-rank framework, have been extensively studied in the literature (Kveton et al., 2015a, b; Li et al., 2016; Wang & Chen, 2018; Vial et al., 2022; Zhong et al., 2021; Liu et al., 2022; Wang et al., 2023, 2024). Offline cascading bandits, on the other hand, focus primarily on reducing bias in learning settings (Joachims, 2002; Wang et al., 2018, 2016; Keane & O’Brien, 2006; Zhang et al., 2023). Unlike these prior works, our study tackles the unbiased setting where data coverage is insufficient. Moreover, we are the first to provide a theoretically guaranteed solution using a CMAB-based approach.
LLM caching is a memory management technique aimed at mitigating memory footprints and access overhead during training and inference. Previous studies have investigated LLM caching at various levels, including attention-level (KV-cache) (Pope et al., 2022; Kwon et al., 2023; Sheng et al., 2023; Bang, 2023), query-level (Gim et al., 2023; Zhu et al., 2023), and model/API-level (Qu et al., 2024; Dai et al., 2024a; Feng et al., 2024). Among these, the closest related work is the LLM cache bandit framework proposed by Zhu et al. (2023). However, their approach is ad hoc, whereas our CMAB-based framework systematically tackles the same problem and achieves improved results in both offline and online settings.
Influence Maximization (IM) was initially formulated as an algorithmic problem by Richardson & Domingos (2002) and has since been studied using greedy approximation algorithms (Kempe et al., 2003b; Chen et al., 2009). The online IM problem has also received significant attention (Wen et al., 2017; Vaswani et al., 2017a; Wu et al., 2019; Vaswani et al., 2017b, 2015; Wang & Chen, 2017). In the offline IM domain, our work aligns closely with the optimization-from-samples (OPS) framework (Balkanski et al., 2015, 2016; Chen et al., 2020, 2021), originally proposed by Balkanski et al. (2015). Specifically, our work falls under the subdomain of optimization-from-structured-samples (OPSS) (Chen et al., 2020, 2021), where samples include detailed diffusion step information $(S_{0},...,S_{V-1})$ instead of only the final influence spread $\sigma(S_{0};G)$ in the standard OPS. Compared to Chen et al. (2021), which selects the best seed set using empirical means, our approach employs a variance-adaptive pessimistic LCB, improving the suboptimal gap under relaxed assumptions.
Appendix B More Justification of Studying Offline CMAB
While online bandits are a natural choice when online data is readily available and inexpensive, many real-world applications restrict access to only offline data as follows, which motivates the study of offline CMAB.
For instance, consider the cascading bandit model in recommendation systems (Kveton et al., 2015a). Online CMAB learning requires a tight feedback loop where the platform (i.e., the learner) updates its recommendation policy after every user interaction. However, in many practical scenarios, such fine-grained online feedback is unavailable as the platform cannot afford to update at such a high frequency. Instead, data is collected in batches (e.g., over a week), logged, and then used to update the policy in a single offline training phase. This workflow aligns precisely with our offline CMAB setting.
Another motivating scenario involves outsourced system design. For example, if OpenAI or Anthropic outsources the design of an LLM caching system, the consultant (i.e., the learner) typically receives only anonymized user logs. They must learn user behavior and design the system purely based on this private offline dataset and cannot reach out for direct interaction with the users, which fits naturally into the offline CMAB framework.
Moreover, our work on CMAB also mirrors the development trajectory in reinforcement learning (RL). RL began with a focus on online learning (Mnih et al., 2013); then, around 2020, concerns over the cost and availability of online interactions led to a growing emphasis on offline RL—learning solely from logged data (Levine et al., 2020). More recently, hybrid approaches (Lee et al., 2022) combining offline pretraining with online fine-tuning have emerged. Similarly, after establishing foundational results in online CMAB, we now focus on the offline setting as a crucial step toward enabling future hybrid CMAB approaches.
We will incorporate this discussion and examples into the final version of the paper.
Appendix C Offline Learning for Influence Maximization with Extension to Node-level Feedback
Influence maximization (IM) is the task of selecting a small number of seed nodes in a social network to maximize the influence spread from these nodes, which has been applied in various important applications such as viral marketing, epidemic control, and political campaigning (Richardson & Domingos, 2002; Kempe et al., 2003b; Chen et al., 2009). IM has been intensively studied over the past two decades under various diffusion models, such as the independent cascade (IC) model (Kempe et al., 2003b), the linear threshold (LT) model (Chen et al., 2010), and the voter model (Narasimhan et al., 2015)], as well as different feedback such as edge-level and node-level feedback models (Chen et al., 2020). For the edge-level feedback model, IM smoothly fits into our framework by viewing each edge as the base arm, which can obtain the theoretical result similar to our previous two applications. In this section, we consider a more realistic yet challenging setting where we can only obverse the node-level feedback, showing that our framework still applies as long as we can construct a high probability lower bound (LCB) for each base arm (edge).
@@ Appendix C Offline Learning for Influence Maximization with Extension to Node-level Feedback
Similar to (Chen et al., 2021), we assume these $n$ influence cascades are generated independently from a seed set distribution $S_{t,0}\sim\mathbb{D}_{\mathcal{S}}$, and given $S_{t,0}$, the cascades are generated according to the IC diffusion process. For each node $v\in\mathcal{V}$, we use $q_{v}=\Pr\left[v\in S_{t,0}\right]$ to denote the probability that the node $v$ is selected by the experimenter in the seed set $S_{t,0}$. We use $p_{G}(\bar{v})\vcentcolon=\Pr\left[v\notin S_{t,1}\right]$ to denote the probability that the node $v$ is not activated in one time step when the graph is $G$. For any two nodes $u,v\in\mathcal{V}$, we use $p_{G}(\bar{v}|u)\vcentcolon=\Pr\left[v\notin S_{t,1}|u\in S_{t,0}\right]$ and $p_{G}(\bar{v}|\bar{u})\vcentcolon=\Pr\left[v\notin S_{t,1}|u\notin S_{t,0}\right]$ to denote the probability that the node $v$ is not activated in one time step conditioned on whether the node $u$ is in the seed set $S_{t,0}$ or not, respectively. We also assume $\mathbb{D}_{\mathcal{S}}$ is a product distribution, i.e., each node $u\in\mathcal{V}$ is selected as a seed node in $S_{t,0}$ independently. Similar to (Chen et al., 2021), we also need an additional 1.
@@
Inspired by (Chen et al., 2021), for each node-level feedback data $t\in[n]$, we only use the seed set $S_{t,0}$ and the active nodes in the first diffusion step $S_{t,1}$ to construct the LCB.
@@
To find out an action $\widehat{S}$ such that $\sigma(\widehat{S};G)\geq(\alpha-\epsilon)\sigma(S^{*};G)$, our algorithm requires that $n\geq\tilde{O}\left(\frac{V^{2}d_{\max}^{2}}{\epsilon^{2}\eta\gamma^{3}}\right)$, which improves the existing result by at least a factor of $\tilde{O}\left(\frac{V^{4}}{k^{2}d^{2}_{\max}\eta}\right)$, owing to our variance-adaptive LCB construction and the tight CMAB-T analysis. We also relax the assumption regarding 1, where we require bounded $q_{u},p(\bar{v})$ only for $(u,v)\in\mathcal{E}(s^{*})$, since we use LCB $\underaccent{\bar}{p}_{uv}$. Chen et al. (2021), instead, needs bounded $q_{u},p(\bar{v})$ for all $(u,v)\in\mathcal{E}$ as they directly use the empirical mean of $p_{uv}$.
Appendix D Proof for the Upper Bound Result